# Physics and Deep Learning in Computational Wave Imaging

### v1

@@ Abstract
Computational wave imaging (CWI) extracts hidden structure and physical properties of a volume of material by analyzing wave signals that traverse that volume. Applications include seismic exploration of the Earth’s subsurface, acoustic imaging and non-destructive testing in material science, and ultrasound computed tomography in medicine. Current approaches for solving CWI problems can be divided into two categories: those rooted in traditional physics, and those based on deep learning. Physics-based methods stand out for their ability to provide high-resolution and quantitatively accurate estimates of acoustic properties within the medium. However, they can be computationally intensive and are susceptible to ill-posedness and nonconvexity typical of CWI problems. Machine learning-based computational methods have recently emerged, offering a different perspective to address these challenges. Diverse scientific communities have independently pursued the integration of deep learning in CWI. This review delves into how contemporary scientific machine-learning (ML) techniques, and deep neural networks in particular, have been harnessed to tackle CWI problems. We present a structured framework that consolidates existing research spanning multiple domains, including computational imaging, wave physics, and data science. This study concludes with important lessons learned from existing ML-based methods and identifies technical hurdles and emerging trends through a systematic analysis of the extensive literature on this topic.
Index Terms:
Computational Wave Imaging, Deep Learning, Wave Physics, Full Waveform Inversion, Ultrasound Computed Tomography, Non-Destructive Testing
@@ I Introduction
Waves are ubiquitous in our world; prominent examples include the propagation of sound through the atmosphere, waves in the ocean, and seismic waves in the earth. Well-established physical models provide detailed descriptions of how these waves are influenced by the materials through which they propagate. Computational wave imaging (CWI) aims to invert these models, and to extract detailed information about these materials, including properties such as density and bulk modulus, using measurements of wave signals (amplitudes, phases, and travel-times) taken on the outer boundaries of these spaces.
@@
This review will focus on the three distinct application areas shown in Fig. 1a: seismic Full-Waveform Inversion (FWI), acoustic and industrial ultrasound imaging (encompassing Non-Destructive Testing, or NDT), and medical Ultrasound Computed Tomography (USCT). Despite the differences in the frequency of the propagating wave (Fig. 1b), and the scale of the imaging volumes, these applications share several common underlying physical processes. Each involves both a forward wave propagation model and an inverse imaging or reconstruction problem, as depicted in Fig. 1a. “Wave Propagation” is described by the governing wave equation, and is generally the more straightforward problem; it often requires the use of numerical methods to simulate waves in complex media, as analytical solutions to wave equations are only attainable in idealized conditions. Conversely, the more challenging “reconstruction” task is to estimate the properties of the intervening medium based on measurements (e.g., of acoustic pressure) taken outside the region of interest. We have observed the great potential of machine learning (ML) to enhance the accuracy and efficiency of solutions to CWI problem in various fields, such as subsurface energy applications [1, 2, 3], material characterization [4, 5, 6], and medical ultrasound imaging [7, 8, 9]. Such improvements to CWI can be impactful, as these techniques are not only tools for problem-solving but also catalysts for scientific discovery and engineering innovation. Indeed, they can even lead to the discovery of new physical insights [10]. Furthermore, we have found that integrating physics into ML models contributes to improved model generalization and robustness, and makes more effective use of limited training data [11, 12]. Our objective is to provide an overview of data science and ML techniques used for CWI challenges, and to underscore the significant advantages of incorporating the fundamental underlying physics into the design of ML models.
@@
A common set of computational challenges emerges from the shared foundation of wave physics across diverse imaging problems. One prominent challenge is the pervasive ill-posedness encountered in CWI problems, arising from the inherently limited data coverage, i.e. attempting to characterize a 3D volume with measurements confined to its 2D surface. To mitigate ill-posedness, conventional physics-based methods introduce techniques such as regularization, but selecting an appropriate regularization term can be a non-trivial task. Additionally, numerical optimization methods for solving CWI problems have their own difficulties: (1) the accuracy of the solution can be highly contingent on the initial guess; and (2) the computational cost can be prohibitive. Traditional physics-based CWI solvers often require many evaluations of the forward model before achieving convergence, presenting substantial computational burden. Furthermore, cycle-skipping is a well-documented issue in CWI [13] that occurs when the timing of the modeled waveform (predicted by the inverted solution) deviates by more than half a cycle from that of the observed waveform. The goal of this paper is to provide a review of contemporary methods in CWI, with a focus on common challenges and methods. In particular, we explore both the opportunities and difficulties inherent in the application of ML to enhance and expedite the solution of CWI problems.
@@
ML has significantly influenced the evolution of CWI techniques. This synergy is particularly evident in the emerging trend of integrating fundamental physics principles with ML algorithms to advance CWI methods. Here we have compiled over 200 papers from three CWI communities, and have observed a clear and consistent upward trend in the volume of publications dedicated to ML within this domain (as shown in Fig. 9 in the Appendix). Furthermore, each of these scientific communities has produced high-quality review papers, a selection of which is presented in Table II of the Appendix. These reviews can be categorized into three primary groups: physics-based methods, machine-learning approaches, and hybrid methods that integrate data-driven techniques with fundamental physics principles.
@@ II Governing Wave Physics
CWI relies on partial differential equations (PDEs) tailored to various wave types and propagation media. In Earth sciences, the acoustic wave equation is commonly used in exploration geophysics [14], while the elastic wave equation is fundamental in earthquake seismology [15]. In medical imaging, the acoustic wave model is frequently employed in ultrasound computed tomography [16], whereas a coupled elastic-acoustic wave model is used for elastography and transcranial ultrasound [17]. This paper predominantly features the acoustic wave equation due to its effective representation of underlying wave physics and its simplicity in formulation. Computational methods developed using the acoustic wave model can naturally extend to more complex wave physics models, including elastic, anisotropic, and visco-acoustic waves (c.f. Appendix A).
@@ II-A Acoustic Wave Equation
The acoustic wave equation describes the propagation of compressional (P) waves through a non-lossy medium with homogeneous density; specifically
$\nabla^{2}p(\bm{r},t)-\frac{1}{V_{P}^{2}(\bm{r})}\frac{\partial^{2}}{\partial t^{2}}p(\bm{r},t)=s(\bm{r},t)\,,$ (1)
where $t$ is time, $\bm{r}$ is spatial location, $\nabla^{2}$ is the Laplacian operator, and $V_{P}(\bm{r})$ is the velocity map (i.e., the speed of the P wave as a function of position $\bm{r}$). Here, $s(\bm{r},t)$ is the source term that specifies the location and time history of the source, and $p(\bm{r},t)$ is the pressure wavefield. The acoustic wave equation represents just one form among many wave equation variants.
@@
Figure 2: A taxonomy of ML methods for CWI, based on two categorizations. Under supervision strategy, we distinguish fully supervised, semi-supervised, self-supervised, and unsupervised. Under learning strategy, we categorize methods as simulations/measurements-driven, physics-aware, parameterized solutions, and plug-and-play priors. We further differentiate these methods by “Data Consistent” (in green) and “Non-Data Consistent” (in blue) based on their adherence to data fidelity consistency. In Appendix B, we offer visualizations illustrating trends in ML methods for CWI, along with a breakdown categorizing over 200 papers into eight method groups. This section also includes a curated table highlighting selected papers from each category.
@@ III-A Forward Modeling
Physics-based methods are broadly implemented in various CWI applications. Generally speaking, these methods are computationally expensive, and require careful regularization. However, they tend to be more generalizable and robust to measurements contaminated with noise than data-driven methods. The forward modeling of CWI is usually posed as
@@
where $\bm{d}$ represents the observed data, $f(\cdot)$ is the forward model, $\bm{m}$ represents the material properties we wish to infer, and $\eta$ represents measurement noise. For acoustic forward modeling, the observed data $\bm{d}$ consists of time traces of the pressure wavefield $p(\bm{r},t)$ at discrete receiver locations $\bm{r}$ (usually on the exterior of the region whose properties we are attempting to probe), while $\bm{m}$ corresponds to the velocity map $V_{P}(\bm{r})$. In fields such as exploration geophysics and medical ultrasound, the source term $s(\bm{r})$ can usually be assumed to be known since the sources are controlled, and the source function is typically directly measurable [18]. When the source term is unknown, such as in earthquake seismology, specialized methods are required, including joint inversion or source-independent inversion [19] to infer the properties of the medium.
@@ III-B Physics-based Wave Imaging
CWI aims to estimate the material property $\bm{m}$ from the observed data $\bm{d}$; that is, it aims to find an inversion operator $g$ so that $\widehat{\bm{m}}=g(\bm{d})$ approximates the actual property. A minimal requirement for this approximate model is that the forward propagation based on this model should approximate the observed data; that is, $f(\widehat{\bm{m}})\approx\bm{d}$. This implies that $f(g(\bm{d}))\approx\bm{d}$ and for this reason, $g$ is informally referred to as the inverse of $f$. But directly inverting $f$ is generally infeasible due to the ill-posed nature of CWI. To address these challenges, CWI is often formulated as the solution to an optimization problem. The unknown material properties are then estimated as the minimizers $\widehat{\bm{m}}$ of a penalized nonlinear least squares problem and can be written as [37, 38, 39]
@@
where $\left\|\bm{d}-f(\bm{m})\right\|_{2}^{2}$ is the data fidelity term with $||\cdot||_{2}$ denoting the $\ell_{2}$ norm, and $r(\cdot)$ is the regularization term with coefficient $\gamma$. Common choices for $r(\bm{m})$ are Tikhonov regularization or total-variation (TV) regularization, where Tikhonov regularization has the form $r(\bm{m})=\|H\bm{m}\|^{2}_{2}$, where $H$ is a high-pass filter, while TV regularization uses the $\ell_{1}$ norm of the spatial gradient [40]. The choice of regularization reflects the user’s bias regarding the physical form of the model of the intervening medium: the Tikhonov term biases toward smooth models, while TV regularization promotes piecewise smooth estimates.
@@ IV-A Machine Learning Methods by Supervision Strategy
  • Self-Supervised. Labels and samples are provided, but they are not associated.
  • Unsupervised. No labels are provided and only samples (waveform data) are available.
Figure 3: The schematic illustration of the network architecture of (a) InversionNet, a fully-supervised method [41], (b) Auto-Linear, a self-supervised method [10] and (c) UPFWI, an unsupervised method [11].
@@ IV-A1 Fully Supervised Methods
Fully supervised methods are among the most straightforward and widely used approaches in solving CWI problems [1, 41, 42, 43]. The basic idea is to obtain a large training set $\bm{\Phi}$ of pairs $(\bm{d}_{i},\bm{m}_{i})$ and to use that data to train a deep neural network that instantiates the inversion operator $g_{\theta}(\cdot)$. Here, we employ the “InversionNet” method, developed by Wu and Lin [41], as an example to illustrate how to solve CWI problems using a fully supervised learning strategy. Mathematically, InversionNet can be expressed as
@@
$\mathrm{s.t.}\;\bm{\theta}^{*}$ $=\operatorname*{arg\,min}_{\bm{\theta}}\sum_{(\bm{d}_{i},\bm{m}_{i})\in\bm{\Phi}}\mathcal{L}(g_{\bm{\theta}}(\bm{d_{i}}),\,\bm{m}_{i})\,,$
where $\bm{m}$ is the predicted velocity map, $\bm{\theta}$ represents the trainable weights in the inversion network, $g_{\bm{\theta}}(\cdot)$ corresponds to the network itself, $\bm{\Phi}$ is the training dataset with paired samples of $(\bm{d}_{i},\bm{m}_{i})$, and the loss function $\mathcal{L}(\cdot,\,\cdot)$ is commonly defined in terms of the $\ell_{1}$ or $\ell_{2}$ norm, comparing the difference of $\bm{m}_{i}$ with $\widehat{\bm{m}}$. As shown in Fig. 3a, InversionNet utilizes an encoder-decoder structure, processing both input waveform data and output velocity maps as images. The labeled samples (i.e., $\bm{d},\bm{m}$ pairs) are typically synthetically generated to reflect realistic medium profiles, often numbering in the tens of thousands. Various methods, such as the geostatistical approaches detailed by Azevedo et al. [44], are used to create geologically significant $\bm{m}_{i}$-images for seismic inversion. Each waveform $\bm{d}$ is computed using the forward operator $f$, as outlined in Eq. (2). While training requires a large number of labeled samples, a fully trained network can quickly evaluate the inverse operator $g_{\bm{\theta}^{*}}(\cdot)$, an advantage in monitoring scenarios requiring frequent assessments. However, inference on out-of-distribution data remains challenging, and the network tends to generate solutions that closely mirror the training data, thus limiting generalization.
In addition to InversionNet, several other works build on the concept of fully supervised learning strategies, employing various backbone networks. Araya-Polo et al. [1] utilize fully connected networks, while Rao et al. [43] develop an encoder-decoder network structure similar to InversionNet to enhance acoustic imaging. Yang and Ma [42] implement a specific type of encoder-decoder network known as U-net [45], which includes skip connections linking the encoder and decoder paths, to facilitate learning. The choice of backbone network architecture can significantly impact the network’s properties. Generally, convolution-based networks are more parameter-efficient due to weight sharing, making them particularly suitable for processing high-dimensional waveform data and velocity maps while reducing the size of the required training dataset. In contrast, fully connected networks, due to their high parameter count, may become impractical for large-scale applications.

### v2

# Survey of Deep Learning and Physics-Based Approaches in Computational Wave Imaging

@@ Abstract
Computational wave imaging (CWI) extracts hidden structure and physical properties of a volume of material by analyzing wave signals that traverse that volume. Applications include seismic exploration of the Earth’s subsurface, acoustic imaging and non-destructive testing in material science, and ultrasound computed tomography in medicine. Current approaches for solving CWI problems can be divided into two categories: those rooted in traditional physics, and those based on deep learning. Physics-based methods stand out for their ability to provide high-resolution and quantitatively accurate estimates of acoustic properties within the medium. However, they can be computationally intensive and are susceptible to ill-posedness and nonconvexity typical of CWI problems. Machine learning-based computational methods have recently emerged, offering a different perspective to address these challenges. Diverse scientific communities have independently pursued the integration of deep learning in CWI. This review discusses how contemporary scientific machine-learning (ML) techniques, and deep neural networks in particular, have been developed to enhance and integrate with traditional physics-based methods for solving CWI problems. We present a structured framework that consolidates existing research spanning multiple domains, including computational imaging, wave physics, and data science. This study concludes with important lessons learned from existing ML-based methods and identifies technical hurdles and emerging trends through a systematic analysis of the extensive literature on this topic.
@@ I Introduction
Waves are ubiquitous in our world; prominent examples include the transmission of sound through the atmosphere, the motion of water on the surface of the ocean, and the traveling of seismic vibrations through the earth. Well-established physical models provide detailed descriptions of how these waves are influenced by the materials through which they propagate. Computational wave imaging (CWI) aims to invert these models, and to extract detailed information about these materials, including properties such as density and bulk modulus, using measurements of wave signals (amplitudes, phases, and travel-times) taken on the outer boundaries of these spaces.
@@
This review will focus on the three distinct application areas shown in Fig. 9: seismic Full-Waveform Inversion (FWI), acoustic and industrial ultrasound imaging (encompassing Non-Destructive Testing, or NDT), and medical Ultrasound Computed Tomography (USCT). Despite the differences in the frequency of the propagating wave (Fig. 1b), and the scale of the imaging volumes, these applications share several common underlying physical processes. Each involves both a forward wave propagation model and an inverse imaging or reconstruction problem, as depicted in Fig. 9. “Wave Propagation” is described by the governing wave equation, and is generally the more straightforward problem; it often requires the use of numerical methods to simulate waves in complex media, as analytical solutions to wave equations are only attainable in idealized conditions. Conversely, the more challenging “reconstruction” task is to estimate the properties of the intervening medium based on measurements (e.g., of acoustic pressure) taken outside the region of interest. We have observed the great potential of machine learning (ML) to enhance the accuracy and efficiency of solutions to CWI problem in various fields, such as subsurface energy applications [1, 2, 3], material characterization [4, 5, 6], and medical ultrasound imaging [7, 8, 9]. Such improvements to CWI can be impactful, as these techniques are not only tools for problem-solving but also catalysts for scientific discovery and engineering innovation. Indeed, they can even lead to the discovery of new physical insights [10]. Furthermore, we have found that integrating physics into ML models contributes to improved model generalization and robustness, and makes more effective use of limited training data [11, 12]. Our objective is to provide an overview of data science and ML techniques used for CWI challenges, and to underscore the significant advantages of incorporating the fundamental underlying physics into the design of ML models.
@@
A common set of computational challenges emerges from the shared foundation of wave physics across diverse imaging problems. One prominent challenge is the pervasive ill-posedness encountered in CWI problems, arising from the inherently limited data coverage; i.e., attempting to characterize a 3D volume with measurements confined to its 2D surface. To mitigate ill-posedness, conventional physics-based methods introduce techniques such as regularization, but selecting an appropriate regularization term can be a non-trivial task. Additionally, numerical optimization methods for solving CWI problems have their own difficulties: (1) the accuracy of the solution can be highly contingent on the initial guess; and (2) the computational cost can be prohibitive. Traditional physics-based CWI solvers often require many evaluations of the forward model before achieving convergence, presenting substantial computational burden. Furthermore, cycle-skipping is a well-documented issue in CWI [13] that occurs when the timing of the modeled waveform (predicted by the inverted solution) deviates by more than half a cycle from that of the observed waveform. The goal of this paper is to provide a review of contemporary methods in CWI, with a focus on common challenges and methods. In particular, we explore both the opportunities and difficulties inherent in the application of ML to enhance and expedite the solution of CWI problems.
@@
ML has significantly influenced the evolution of CWI techniques. This synergy is particularly evident in the emerging trend of integrating fundamental physics principles with ML algorithms to advance CWI methods. Here we have compiled over 200 papers from three CWI communities, and have observed a clear and consistent upward trend in the volume of publications dedicated to ML within this domain (as shown in Fig. 15 in the Appendix). Furthermore, each of these scientific communities has produced high-quality review papers, a selection of which is presented in Table B of the Appendix. These reviews can be categorized into three primary groups: physics-based methods, machine-learning approaches, and hybrid methods that integrate data-driven techniques with fundamental physics principles.
@@ II Governing Wave Physics
Wave-based imaging relies on partial differential equations (PDEs) to describe the propagation of mechanical waves through physical media. The choice of wave model varies across application domains and is dictated by the properties of the medium and the nature of the wave. In Earth sciences, the acoustic wave equation is commonly used in seismic exploration [14], while elastic models are essential for simulating earthquake processes [15]. In medical imaging, acoustic wave models are standard for ultrasound computed tomography [16], with elastic or viscoelastic models employed in applications such as elastography and transcranial ultrasound [17]. In nondestructive testing (NDT), both acoustic and elastic wave equations are used, depending on the geometry and material complexity of the structures being interrogated [6].
@@ II-A Acoustic Wave Equation
The acoustic wave equation models the propagation of compressional (P) waves through fluid-like media, where shear stresses are negligible. It is widely used in seismic exploration [14], ultrasound imaging [16], and NDT [6] scenarios involving air or water. This model also serves as the default choice in many learning-based CWI frameworks due to its mathematical simplicity and computational efficiency.
The constant density form of the acoustic wave equation is:
$\nabla^{2}p(\bm{r},t)-\frac{1}{V_{P}^{2}(\bm{r})}\frac{\partial^{2}p(\bm{r},t)}{\partial t^{2}}=s(\bm{r},t),$ (1)
where $p(\bm{r},t)$ denotes the pressure field, $V_{P}(\bm{r})$ is the P-wave velocity, $\bm{r}$ represents spatial coordinates, and $s(\bm{r},t)$ specifies the spatiotemporal source. This form assumes uniform density, simplifying numerical treatment but neglecting key physical effects in heterogeneous media.
The variable-density acoustic wave equation introduces spatial variability in density $\rho(\bm{r})$ [18, 19]:
$\nabla\cdot\left(\frac{1}{\rho(\bm{r})}\nabla p\right)-\frac{1}{\rho(\bm{r})V_{P}^{2}(\bm{r})}\frac{\partial^{2}p}{\partial t^{2}}=s(\bm{r},t).$ (2)
This model offers a more accurate description of wave propagation in complex environments such as sedimentary basins or soft tissues. While computationally more demanding, it remains widely used due to its flexibility and compatibility with both inversion and learning-based algorithms.
Overall, the acoustic wave equation plays a foundational role in CWI. It supports synthetic data generation, enables physics-informed learning, and offers a physically grounded framework for designing loss functions and priors in modern deep learning pipelines. Unless otherwise noted, the remainder of this paper adopts the constant density formulation without loss of generality.
II-B Extensions to Complex Wave Physics
Although the acoustic wave equation provides a foundational model widely used across many CWI applications, real-world propagation environments often involve more complex physical phenomena. These complexities arise from material anisotropy, energy dissipation, and the presence of both compressional and shear wave modes. To address these effects, various extensions to the acoustic model have been developed and are employed in geophysics, medical imaging, and nondestructive testing [20, 21, 22].
Elastic Wave Equation.  In solid media, wave propagation involves both compressional (P) and shear (S) waves. The elastic wave equation models this behavior by jointly solving for velocity and stress tensors, using Lamé parameters to describe the material’s elastic properties. It is essential in applications such as earthquake seismology and materials characterization [15].
Viscoacoustic Wave Equation.  To model attenuation in lossy or dissipative media, viscoacoustic formulations introduce memory variables and relaxation mechanisms that simulate the frequency-dependent absorption of wave energy [23]. These models are critical in biomedical and seismic contexts where wave dispersion and amplitude decay significantly impact image resolution and accuracy.
Anisotropic Wave Equation.  Many materials exhibit directionally dependent wave speeds due to internal structure or stress fields. Anisotropic wave equations, including simplified pseudo-acoustic forms such as Vertical Transverse Isotropy (VTI), enable modeling of wave propagation in these directionally varying environments [21, 24]. Such models are widely used in layered earth models and anisotropic composites.
Other Variants.  Additional extensions include poroelastic models for fluid-solid interactions and coupled elastic-acoustic systems [25]. These formulations offer greater fidelity but also introduce increased complexity in both numerical simulation and learning-based inversion.
While this review focuses on the acoustic wave equation for clarity and broad applicability, many of the learning methods described in the following sections can be extended to handle these more complex wave physics. Readers are referred to Appendix A for detailed mathematical formulations of these advanced wave equations.
@@ II-B Extensions to Complex Wave Physics
Figure 2: A taxonomy of ML methods for CWI, based on two categorizations. Under supervision strategy, we distinguish fully supervised, semi-supervised, self-supervised, and unsupervised/unlabeled. Under learning strategy, we categorize methods as simulations/measurements-driven, physics-aware, parameterized solutions, and plug-and-play priors. We further differentiate these methods by “Data Consistent” (in green) and “Non-Data Consistent” (in blue) based on their adherence to data fidelity consistency. In Appendix B, we offer visualizations illustrating trends in ML methods for CWI, along with a breakdown categorizing over 200 papers into eight method groups. This section also includes a curated table highlighting selected papers from each category.
@@ III-A Forward Modeling
Physics-based methods are broadly implemented in various CWI applications. Generally speaking, these methods are computationally expensive, and require careful regularization. However, they tend to be more generalizable and robust to measurements contaminated with noise than data-driven methods. The forward modeling of CWI can be written
@@
where $\bm{d}$ represents the observed data, $f(\cdot)$ is the forward model, $\bm{m}$ represents the material properties we wish to infer, and $\eta$ represents measurement noise.
Particularly for acoustic wave modeling, which is the focus of this manuscript, the observed data $\bm{d}$ typically consists of time traces of the pressure wavefield $p(\bm{r},t)$ recorded at discrete receiver locations $\bm{r}$, while $\bm{m}$ corresponds to the spatially varying velocity map $V_{P}(\bm{r})$. In many practical settings, the forward model also depends on additional physical or algorithmic nuisance parameters—quantities that influence the data but are not the primary inversion targets. Common examples include the source location, source time function, boundary conditions, or sensor calibration parameters. While not the primary focus of CWI, these factors must be carefully handled to ensure accurate reconstruction.
@@ III-B Physics-based Wave Imaging
Computatioanl wave imaging aims to estimate the material property $\bm{m}$ from the observed data $\bm{d}$; that is, it aims to find an inversion operator $g$ so that $\widehat{\bm{m}}=g(\bm{d})$ approximates the actual property. A minimal requirement for this approximate model is that the forward propagation based on this model should approximate the observed data; that is, $f(\widehat{\bm{m}})\approx\bm{d}$. This implies that $f(g(\bm{d}))\approx\bm{d}$ and for this reason, $g$ is informally referred to as the inverse of $f$. But directly inverting $f$ is generally infeasible due to the ill-posed nature of CWI. To address these challenges, CWI is often formulated as the solution to an optimization problem. The unknown material properties are then estimated as the minimizers $\widehat{\bm{m}}$ of a penalized nonlinear least squares problem and can be written
@@
where the first term enforces data fidelity and the second term, weighted by $\gamma$, incorporates prior information with a regularization function $r(\bm{m})$, which encodes assumptions about the underlying medium [43, 44, 45, 46]. Classical choices include Tikhonov regularization, $r(\bm{m})=\|H\bm{m}\|_{2}^{2}$, which promotes smoothness via a high-pass filter $H$, and total variation (TV) regularization, $r(\bm{m})=\|\nabla\bm{m}\|_{1}$, which encourages piecewise smooth structures and preserves edges[47].
Beyond these, many advanced regularization strategies have been developed to better adapt to specific medium structures and improve reconstruction quality. For instance, modified TV formulations have been used to enhance the delineation of high-contrast interfaces [45]; adaptive schemes such as the proximal Newton–ADMM framework with pretrained denoisers allow dynamic balancing between smooth and blocky regions [48]; and constraint-based approaches like salt flooding techniques have proven effective in complex subsurface environments [49]. Recently, hybrid strategies that decompose the model into smooth and blocky components, such as the adaptive Tikhonov–TV regularization method [50], demonstrated strong robustness across varying realistic settings.
@@ IV-A Machine Learning Methods by Supervision Strategy
  • Self-Supervised. Labels are not provided, but label-like supervision signals (pseudo labels) are generated from the samples.
  • Unsupervised/Unlabeled.  Training is performed using only input samples, without explicit labels. This category includes both data-driven methods that learn latent structure (e.g., autoencoders, contrastive learning), as well as learning approaches that infer velocity parameters by enforcing consistency with simulated or observed measurements.
Figure 3: Network architecture of InversionNet, a fully-supervised method [51].
@@ IV-A1 Fully Supervised Methods
Fully supervised methods are among the most straightforward and widely used approaches in solving CWI problems [1, 51, 52, 53]. The basic idea is to obtain a large training set $\bm{\Phi}$ of pairs $(\bm{d}_{i},\bm{m}_{i})$ and to use that data to train a deep neural network that instantiates the inversion operator $g_{\theta}(\cdot)$. Here, we employ the “InversionNet” method, developed by [51], as an example to illustrate how to solve CWI problems using a fully supervised learning strategy. Mathematically, InversionNet can be expressed as
@@
$\mathrm{where}\;\bm{\theta}^{*}$ $=\operatorname*{arg\,min}_{\bm{\theta}}\sum_{(\bm{d}_{i},\bm{m}_{i})\in\bm{\Phi}}\mathcal{L}(g_{\bm{\theta}}(\bm{d_{i}}),\,\bm{m}_{i})\,$,
where $\bm{m}$ is the predicted velocity map, $\bm{\theta}$ represents the trainable weights in the inversion network, $g_{\bm{\theta}}(\cdot)$ corresponds to the network itself, $\bm{\Phi}$ is the training dataset with paired samples of $(\bm{d}_{i},\bm{m}_{i})$, and the loss function $\mathcal{L}(\cdot,\,\cdot)$ is commonly defined in terms of the $\ell_{1}$ or $\ell_{2}$ norm, comparing the difference of $\bm{m}_{i}$ with $\widehat{\bm{m}}$. As shown in Fig. 3, InversionNet utilizes an encoder-decoder structure, processing both input waveform data and output velocity maps as images. The labeled samples (i.e., $\bm{d},\bm{m}$ pairs) are typically synthetically generated to reflect realistic medium profiles, often numbering in the tens of thousands. Various methods, such as the geostatistical approaches detailed by [54], are used to create geologically significant $\bm{m}_{i}$-images for seismic inversion. Each waveform $\bm{d}$ is computed using the forward operator $f$, as outlined in Eq. (3). While training requires a large number of labeled samples, a fully trained network can quickly evaluate the inverse operator $g_{\bm{\theta}^{*}}(\cdot)$, an advantage in monitoring scenarios requiring frequent assessments. However, inference on out-of-distribution data remains challenging, and the network tends to generate solutions that closely mirror the training data, thus limiting generalization.
In addition to InversionNet, several other works build on the concept of fully supervised learning strategies, employing various backbone networks. [1] utilizes fully connected networks, while [53] develops an encoder-decoder network structure similar to InversionNet to enhance acoustic imaging. [52] implements a specific type of encoder-decoder network known as U-net [55], which includes skip connections linking the encoder and decoder paths, to facilitate learning. The choice of backbone network architecture can significantly impact the network’s properties. Generally, convolution-based networks are more parameter-efficient due to weight sharing, making them particularly suitable for processing high-dimensional waveform data and velocity maps while reducing the size of the required training dataset. In contrast, fully connected networks, due to their high parameter count, may become impractical for large-scale applications.